OpenStudy (mayaal):

giving a medal! find all solutions in the interval(0,2pi) sin^x-cos^x=0

OpenStudy (mathmale):

Need clarification. What do you mean by " sin^x "?

OpenStudy (mayaal):

oops,sorry:) it is: sin^2x-cos^2x=0

OpenStudy (mathmale):

or even better, (sin x)^2 - (cos x)^2 = 0. What have you thought of doing so far?

OpenStudy (mayaal):

i tried using the power reducing formulas and substituting them in this equation.

OpenStudy (mayaal):

am i on the right track?

OpenStudy (mathmale):

You might find it easier to use a substitution for either (sin x)^2 or (cos x)^2. Which trig identity would apply nicely here?

OpenStudy (mayaal):

pythagorean trig identity? sin^2x+cos^2x=1

OpenStudy (mayaal):

@mathmale ?

OpenStudy (mathmale):

that's be neat. from that identity you could obtain (cos x)^2 = 1 - (sin x)^2. Want to try that? Doing so will eliminate the (cos x)^2 from your equation and leave you with (cos x)^2 alone.

OpenStudy (mayaal):

ok: sin^2x-(1-sin^2x)=0 sin^2x-1+sin^2x=0 2sin^2x-1=0

OpenStudy (mathmale):

Yes, and then 2 (sin x)^2 = 1, and (sin x)^2 = 1/2. What next? There will be more than one solution on the given interval.

OpenStudy (mayaal):

so,the nest step is: sin x=square root of 1/2 is this right?

OpenStudy (mathmale):

Yes, except that you need to write sin x = plus or minus 1/Sqrt(2).

OpenStudy (mayaal):

so,what r the solutions?

OpenStudy (mayaal):

x=pi/4,3pi/4,5pi/4,7pi/4?

OpenStudy (mayaal):

does anyone want to help me?

OpenStudy (mayaal):

@mathmate ?

OpenStudy (mayaal):

anybody there??????

OpenStudy (anonymous):

Your answer is correct!

OpenStudy (mayaal):

Oh,thank god.Thankyou very much:)

OpenStudy (anonymous):

You are welcome :)